For a binomial process, the probability of success is 40 percent and the number of trails is 5. Find the variance.

5.0

1.2

2.0

1.1
Answer #1

Solution :

Given that,

p = 0.40

q = 1 - p = 1 - 0.40 = 0.60

n = 5

Using binomial distribution,

variance = 2  = n * p * q = 5 * 0.40 * 0.60

variance = 2  = 1.2